I need help with this equation, Janes mom is 32 yrs old, when jane was born, now janes mom is 3 times old as jane. how old is jane?
Answer by stanbon(75887)

Janes mom was 32 yrs old, when jane was born,
Now jane's mom is 3 times old as jane. how old is jane?
----
Equations:
m = 3j
-----
m-32 = j
===============
Substitute for "j" and solve for "m":
m = 3(m-32)
m = 3m-96
2m = 96
m = 48 (mom's age now)
----
Solve for "j":
j = m-32
j = 48-32 = 16 (jane's age now)
==================================
Cheers,
Stan H.